# thanks: deceasedthanks: deceased

### v1

@@ IV.2 $m_{\pi Y}$ distributions
To investigate the role of $Y^{*}$-resonances, if the events coupling to the $Y^{*}$-poles, we plotted the invariant-mass distribution of the $[\pi Y]$ sub-system, specifically $d\sigma/dm_{[\pi Y]}$, as shown in Fig. 8. For the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel, we have excluded events originating from the direct-2NA process by removing events with $p_{p_{\rm miss}}^{\rm lab}<0.3~{}{\rm GeV}/c$. This criterion facilitates a more straightforward comparison with other channels.
The figures illustrate that the overall shapes of the $m_{\pi Y}$ distributions are similar, exhibiting a clear two-body decay of $Y^{*}\to\pi Y$ over a broad non-resonant background having similar yields. Regarding the $[\pi^{\pm}\Lambda]$ distributions, the positions of the peaks are consistent with those of $\Sigma(1385)^{\pm}$. This consistency can be attributed to the fixed isospin of $\pi^{\pm}\Lambda$, resulting in a natural coupling to $\Sigma(1385)^{\pm}$. In contrast, the observed peaks in the $[\pi^{\mp}\Sigma^{\pm}]$ distributions are slightly higher than $1.4~{}{\rm GeV}/c^{2}$. Hence, it can be inferred that these peaks primarily arise from the coupling to the $\Lambda(1405)$ pole. In this channel, the isospin can be either zero or one. However, the contribution from $\Sigma(1385)^{0}\to\pi^{\mp}\Sigma^{\pm}$ is expected to be negligible. This is because the decay branching ratio of $\Sigma(1385)$, $Br(\pi\Sigma)/Br(\pi\Lambda)$, is approximately 13%, and if the isospin symmetry holds between $\Sigma(1385)^{0}$ and $\Sigma(1385)^{\pm}$, their production yields should be of the same order. No higher mass $Y^{*}$ resonances were observed in the $[\pi Y]$ distributions.
In addition to the study of the $[\pi Y]$ subsystem, we investigated the $[\pi N]$ subsystem to identify if there are any spectral anomalies associated with $\Delta$ or $N^{*}$ resonances. The reactions under consideration were:
@@
The resulting $d\sigma/dm_{\pi N}$ spectra are presented in Fig. 9. In contrast to the $m_{\pi Y}$ distributions, the $m_{\pi N}$ distributions did not exhibit any clear peak or structure.
The obtained results indicate that the dominant processes involve a backward low-momentum $\bar{K}$ in the intermediate state of the reaction via the nucleon knockout reaction $K^{-}N\to\bar{K}N^{\prime}$. This $\bar{K}$ could efficiently couple with $\Sigma(1385)$ or $\Lambda(1405)$ via $\bar{K}+NN\to Y^{*}N\to\pi Y\!N$.
@@ V Discussion
As previously described, the overall structures of the mesonic $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions (see Fig. 6) are similar to the non-mesonic $(m_{[\Lambda p]},q_{[\Lambda p]})$ distribution [2], assuming the phase-space suppression of the $\bar{K}\!N\!N$ peak in mesonic decay channels. The $m_{[\pi Y]}$ spectra presented in Fig. 8 consist of a prominent peak associated with the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, as well as a broad non-resonant contribution.
Therefore, we examined if the event distribution of $\pi Y\!N\!N$ final state can be described essentially in the same model functions applied for $\Lambda pn$ final state, which is composed of $\bar{K}\!N\!N$ formation and ${\rm QF}_{\bar{K}-{\rm abs}}$ as it is described in Ref. [2]. Thus, we assumed that the same event distribution for the intermediate $(\bar{K}+NN)$ system is formed over $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]})$ kinematical plane as it is the case for $\Lambda pn$ final state, and the intermediate $(\bar{K}+NN)$ decays to $\pi Y\!N$. To allow the $\pi Y$ subsystem to couple to the $Y^{*}$-poles in the decay process, we modified the model fitting function $F$ and the phase space accordingly. Therefore, the present model fitting function becomes as follows:
@@
where $\rho$ represents the four-body phase-space for $\pi Y\!N\!N$, $\varepsilon$ denotes the acceptance, $f_{BW}$ is a Breit-Wigner function used to describe the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, $A_{Y^{*}}$ is the coupling strength to the $Y^{*}$ pole, and $f_{j}$ and $A_{j}$ represent the spectral function and strength of each physical process ($j=\bar{K}\!N\!N$ and ${\rm QF}_{\bar{K}-{\rm abs}}$), respectively. The parameters $m_{Y^{*}}$ and $\Gamma_{Y^{*}}$ of $f_{BW}$ were fixed to the values provided by the Particle Data Group (PDG) for the case of $\Sigma(1385)$ [17], while we allowed to vary these values for the case of $\Lambda(1405)$. In the fitting procedure, we ignored any interference between reaction processes. A resolution function, which smears $f_{i}$ and $f_{BW}$, was evaluated by the Monte Carlo simulation and convoluted to Eq.(13).
The $\bar{K}\!N\!N$ spectral function $f_{K}$ is defined as follows:
@@
where $M_{K}$, $\Gamma_{K}$, $Q_{K}$, and $A_{0}^{K}$ are parameters. The ${\rm QF}_{\bar{K}-{\rm abs}}$ spectral function $f_{F}$ is a Gaussian distribution along the $M_{F}(q_{\pi Y\!N})$ (Eq.(10)) given by:
@@
To illustrate how the phase space difference can modify spectral functions, we present Fig. 10, which shows two-dimensional model fitting functions for the $\Lambda p$ and $[\pi Y\!N]$ spectra. The model functions for the $\bar{K}\!N\!N$ production process in each final state are shown in Fig. 10-(a1-3). In the $[\Lambda p]+n$ final state (Fig. 10-(a1)), the reaction threshold ($m_{\Lambda}+m_{N}$) is significantly smaller than $M_{K}$ in Eq.(14), resulting in a nearly flat density-of-state for $[\Lambda p]$ around the $\bar{K}\!N\!N$ peak. On the other hand, in the $[\pi\Lambda N]+N$ (Fig. 10-a2) and $[\pi^{\mp}\Sigma^{\pm}p]+n$ final states (Fig. 10-a3), the reaction threshold ($m_{\pi}+m_{Y}+m_{N}$) approach to $M_{K}$, resulting in substantial suppression of the spectrum due to the diminishing density-of-state at the lighter mass side of the peak structure, so that the maximum of the spectrum appears to shift towards the heavier side compared to $M_{K}$ value. Regarding the ${\rm QF}_{\bar{K}-{\rm abs}}$ process, the spectral modifications in the $[\pi Y\!N]+N$ channels are relatively minor compared to those observed in the $\bar{K}\!N\!N$ case, as illustrated in Fig. 10-(b1-3). This minor modification in the ${\rm QF}_{\bar{K}-{\rm abs}}$ process is primarily due to the energy region being further separated from the reaction threshold.
@@
Figure 10: Fit functions for (a1, a2, a3) the $\bar{K}\!N\!N$ production process and (b1, b2, b3) the quasi-free kaon absorption on the two-nucleon $QF_{\bar{K}-abs}$ process. (a1) and (b1) display spectral functions using the three-body phase space of $\Lambda pn$. (a2) and (b2) correspond to spectral functions using the four-body phase space of $\pi\Lambda NN$. (a3) and (b3) represent spectral functions using the four-body phase space of $\pi\Sigma pn$. The $z$ scale is logarithmic. The black, blue, and red dotted lines correspond to the same lines as in Fig. 6.
@@
This analysis assumes that the detected particles, $\pi Y\!N$, originate from the $(\bar{K}+NN)$ intermediate state. However, it is possible to detect knocked-out $N^{\prime}$ in the primary $K^{-}N\to\bar{K}N^{\prime}$ reaction without detecting $N$ from the intermediate state. In such cases, the $(m,q)$ values of the $[\pi Y]N^{\prime}$ system (rather than $[\pi Y\!N]$) are distributed widely across the kinematically allowed region. To incorporate this scenario, we generated expected spectral functions through simulations for each relevant process and summed them up to account for the background, ${\rm BG}(m,~{}q,~{}m_{\pi Y})$. There is background contamination resulting from the misidentification of the final state, as described in Sec. III.2. In the present fitting, we omitted this background but considered its effect on evaluating the cross-section a systematic error.
@@
Due to the limited statistics over the limited acceptance on three-dimensional kinematical space of $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]},\,m_{[\pi Y]})$, a multi-dimensional fitting nor the application of the coupled channel based sophisticated fitting function (to account for the threshold effects) are both not feasible in this analysis. Instead, we performed a simultaneous fitting on three one-dimensional spectra along the $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$ axes. The two charged modes, $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$, were fitted together by using common values of $M_{\Lambda(1405)}$ and $\Gamma_{\Lambda(1405)}$. The fitting results are shown in Fig. 11, and the obtained parameters are summarized in Table 1. The chi-squares and the number of degrees of freedom for $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ were $\chi^{2}/NDF=275.5/118$, $76.3/85$, and $114.9/136$, respectively. The figures demonstrate that the proposed model functions provide a good overall data description.
@@
Figure 11: The fitted distributions and the results of the fitting for the following channels: (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. “BS”, “QF”, and “BG” represent the $\bar{K}\!N\!N$ production, the ${\rm QF}_{\bar{K}-{\rm abs}}$, and the background processes, respectively.
Table 1: Parameters obtained by the fit.
@@
In Fig. 11-(c) and (d), it can be observed that the $\Lambda(1405)$ contribution in the $m_{\pi^{\mp}\Sigma^{\pm}}$ distributions is well reproduced by a simple Breit-Wigner function. The obtained mass and width of $\Lambda(1405)$ are $1432\pm 5~{}{\rm MeV}/c^{2}$ and $49\pm 3~{}{\rm MeV}/c^{2}$, respectively. These values are close to the pole position deduced from the $d(K^{-},n)\pi\Sigma$ reactions, $(m,\Gamma)=(1418~{}{\rm MeV}/c^{2},52~{}{\rm MeV}/c^{2})$[18], as well as the theoretically predicted pole positions based on chiral unitary approaches, $(1421-1434~{}{\rm MeV}/c^{2},20-52~{}{\rm MeV}/c^{2})$ [19, 20, 21, 22, 23, 24].
@@
We examined the validity of the global fit using $q_{[\pi Y\!N]}$-sliced mass spectra, as shown in Figs. 12 and 13. The model functions successfully reproduced the global structure of each sliced distribution except for the $m_{[\pi Y]}$ spectra near the $\bar{K}\!N$ threshold in the low momentum transfer region of $q<0.3~{}{\rm GeV}/c$ as shown in Fig.13. This discrepancy may come from the neglected channel coupling effect with the $\bar{K}\!N$ channel, i.e., the $\bar{K}N$ threshold cusp in the $I_{\bar{K}N}=1$ channel recently reported by the Belle collaboration [25].
@@
Figure 12: Differential cross sections $d\sigma/dm_{\pi Y\!N}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fit result obtained from the analysis. In panel (a), the gray hatched box highlights a region outside the fitting range to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N\!N}$ threshold.
@@
Figure 13: Differential cross sections $d\sigma/dm_{\pi Y}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fit result obtained from the analysis. In the case of (a), requirements of $m_{\pi^{-}\Lambda p}<2.6~{}{\rm GeV}/c^{2}$ and $p_{\rm miss}>0.3~{}{\rm GeV}/c$ are imposed to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N}$ threshold.
@@
The total cross sections of $\bar{K}\!N\!N$ decaying into each mesonic decay channel, $\sigma_{\bar{K}\!N\!N}^{tot}\,Br(\pi Y\!N)$, were evaluated by integrating the $\bar{K}\!N\!N$ fitting function over the entire kinematical region including above the $\bar{K}$ binding threshold, $m_{\bar{K}}+2m_{N}$ as summarised in Tab. 2. We also tabulated the branch below the $m_{\bar{K}}+2m_{N}$ threshold to exclude the energy region where the spectral shape of the fitting functions would become less reliable due to threshold effects. The values of non-mesonic decay modes, $\bar{K}\!N\!N_{I_{3}=+1/2}\to\Lambda p$ and $\Sigma^{0}p$, were taken from Ref. [2]. The values of the $\bar{K}\!N\!N\to\Sigma^{+}n$ and $\bar{K}\!N\!N\to\pi^{0}\Lambda p$ channels were evaluated as $Br(\Sigma^{+}n)=2Br(\Sigma^{0}p)$ and $Br(\pi^{0}\Lambda p)=1/2Br(\pi^{+}\Lambda n)$, respectively, based on the isospin symmetry. The total non-mesonic decay branch was then evaluated as the sum of the three non-mesonic channels, the errors of which are the linear sum of those three channels since the errors are correlated.
@@
We obtained that the ratio of non-mesonic to mesonic decay branches is $\mathcal{O}(10)$ if we integrate all mass regions. The mesonic decay branch is still larger than the non-mesonic one, even if we integrate below the $\bar{K}$ binding threshold. The result suggests that the mesonic decay mode is the dominant decay branch of the $\bar{K}\!N\!N$, as expected from theoretical considerations.
@@
The decay branches $Br(\pi^{+}\Lambda n)$ and $Br(\pi^{\mp}\Sigma^{\pm}p)$ were similar in magnitude in both integration ranges. The result suggests that the mesonic decay modes coupled to the $I_{\bar{K}\!N}=1$ channel play a significant role in the $\bar{K}\!N\!N$ decay process. Furthermore, this could explain why the decay width of the $\bar{K}\!N\!N$ is much broader than that of the $\Lambda(1405)$ resonance, as the mesonic decay modes with $I_{\bar{K}\!N}=1$ contribute to the broadening of the $\bar{K}\!N\!N$ decay width.
@@
Table 2: Cross sections of each decay channel of $\bar{K}\!N\!N$. The first and second errors are statistical and systematic ones, respectively. The $\bar{K}$ binding thresholds for $I_{3}=\pm 1/2$ were considered as $m_{K^{-}}+2m_{p}$ and $m_{\bar{K}^{0}}+2m_{n}$, respectively. We estimated decay branches of $\Sigma^{+}n$ and $\pi^{0}\Lambda p$ channels by assuming the isospin symmetry.
@@
Decay channel all-region below the $\bar{K}$ binding threshold

### v2

@@ IV.2 $m_{\pi Y}$ distributions
To investigate the role of $Y^{*}$ resonances in the events coupling to the $Y^{*}$ poles, we plotted the invariant-mass distribution of the $[\pi Y]$ subsystem—specifically $d\sigma/dm_{[\pi Y]}$ (Fig. 8). For the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel, we have excluded events originating from the direct-2NA process by removing events with $p_{p_{\rm miss}}^{\rm lab}<0.3~{}{\rm GeV}/c$. This criterion facilitates a more straightforward comparison with other channels.
The figures illustrate that the overall shapes of the $m_{\pi Y}$ distributions are similar, exhibiting a clear two-body decay of $Y^{*}\to\pi Y$ over a broad nonresonant background having similar yields. Regarding the $[\pi^{\pm}\Lambda]$ distributions, the positions of the peaks are consistent with those of $\Sigma(1385)^{\pm}$. This consistency is attributable to the fixed isospin of $\pi^{\pm}\Lambda$, resulting in a natural coupling to $\Sigma(1385)^{\pm}$. By contrast, the peaks observed in the $[\pi^{\mp}\Sigma^{\pm}]$ distributions are slightly greater than $1.4~{}{\rm GeV}/c^{2}$. Hence, we infer that these peaks primarily arise from the coupling to the $\Lambda(1405)$ pole. In this channel, the isospin can be either zero or one. However, the contribution from $\Sigma(1385)^{0}\to\pi^{\mp}\Sigma^{\pm}$ is expected to be negligible because the decay branching ratio of $\Sigma(1385)$, $Br(\pi\Sigma)/Br(\pi\Lambda)$, is approximately 13%; if the isospin symmetry holds between $\Sigma(1385)^{0}$ and $\Sigma(1385)^{\pm}$, their production yields should be of the same order. No higher-mass $Y^{*}$ resonances were observed in the $[\pi Y]$ distributions.
In addition to studying the $[\pi Y]$ subsystem, we investigated the $[\pi N]$ subsystem to determine if there are any spectral anomalies associated with $\Delta$ or $N^{*}$ resonances. The considered reactions are
@@
The resultant $d\sigma/dm_{\pi N}$ spectra are presented in Fig. 9. In contrast to the $m_{\pi Y}$ distributions, the $m_{\pi N}$ distributions did not show a clear peak or structure.
The results indicate that the dominant processes involve a backward low-momentum $\bar{K}$ in the intermediate state of the reaction via the nucleon knockout reaction $K^{-}N\to\bar{K}N^{\prime}$. This $\bar{K}$ can efficiently couple with $\Sigma(1385)$ or $\Lambda(1405)$ via $\bar{K}+NN\to Y^{*}N\to\pi Y\!N$.
@@ V Discussion
As previously described, the overall structures of the mesonic $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions (see Fig. 6) are similar to the non-mesonic $(m_{[\Lambda p]},q_{[\Lambda p]})$ distribution [2], assuming phase-space suppression of the $\bar{K}\!N\!N$ peak in mesonic decay channels. The $m_{[\pi Y]}$ spectra presented in Fig. 8 consist of a prominent peak associated with the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, as well as a broad non-resonant contribution.
Therefore, we examined whether the event distribution of the $\pi Y\!N\!N$ final state can be described essentially by the same model functions applied for the $\Lambda pn$ final state, which is composed of $\bar{K}\!N\!N$ formation and ${\rm QF}_{\bar{K}-{\rm abs}}$, as described in Ref. [2]. Thus, we assumed that the same event distribution for the intermediate $(\bar{K}+NN)$ system is formed over the $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]})$ kinematical plane, as in the case for the $\Lambda pn$ final state, and that the intermediate $(\bar{K}+NN)$ decays to $\pi Y\!N$. To allow the $\pi Y$ subsystem to couple to the $Y^{*}$ poles in the decay process, we modified the model fitting function $F$ and the phase space accordingly. Therefore, the present model fitting function becomes
@@
where $\rho$ represents the four-body phase-space for $\pi Y\!N\!N$, $\varepsilon$ denotes the acceptance, $f_{BW}$ is a Breit–Wigner function used to describe the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, $A_{Y^{*}}$ is the coupling strength to the $Y^{*}$ pole, and $f_{j}$ and $A_{j}$ represent the spectral function and strength of each physical process ($j=\bar{K}\!N\!N$ and ${\rm QF}_{\bar{K}-{\rm abs}}$), respectively. The parameters $m_{Y^{*}}$ and $\Gamma_{Y^{*}}$ of $f_{BW}$ were fixed to the values provided by the Particle Data Group (PDG) in the case of $\Sigma(1385)$ [17], whereas we allowed these values to vary in the case of $\Lambda(1405)$. In the fitting procedure, we ignored any interference between reaction processes. A resolution function, which smears $f_{i}$ and $f_{BW}$, was evaluated by the Monte Carlo simulation and convoluted to Eq. (13).
The $\bar{K}\!N\!N$ spectral function $f_{K}$ is defined as
@@
where $M_{K}$, $\Gamma_{K}$, $Q_{K}$, and $A_{0}^{K}$ are parameters. The ${\rm QF}_{\bar{K}-{\rm abs}}$ spectral function $f_{F}$ is a Gaussian distribution along the $M_{F}(q_{\pi Y\!N})$ (Eq. (10)) given by
@@
To illustrate how the phase space difference can modify spectral functions, we present Fig. 10, which shows two-dimensional model fitting functions for the $\Lambda p$ and $[\pi Y\!N]$ spectra. The model functions for the $\bar{K}\!N\!N$ production process in each final state are shown in Fig. 10(a1–a3). In the $[\Lambda p]+n$ final state (Fig. 10(a1)), the reaction threshold ($m_{\Lambda}+m_{N}$) is substantially smaller than $M_{K}$ in Eq. (14), resulting in a nearly flat density of states for $[\Lambda p]$ around the $\bar{K}\!N\!N$ peak. However, in the $[\pi\Lambda N]+N$ (Fig. 10(a2)) and $[\pi^{\mp}\Sigma^{\pm}p]+n$ final states (Fig. 10(a3)), the reaction threshold ($m_{\pi}+m_{Y}+m_{N}$) approaches $M_{K}$, resulting in substantial suppression of the spectrum because of the diminishing density of states at the lower-mass side of the peak structure; consequently, the maximum of the spectrum appears to shift toward the higher-mass side compared with the $M_{K}$ value. Regarding the ${\rm QF}_{\bar{K}-{\rm abs}}$ process, the spectral modifications in the $[\pi Y\!N]+N$ channels are relatively minor compared with those observed in the $\bar{K}\!N\!N$ case (Fig. 10(b1–b3)). This minor modification in the ${\rm QF}_{\bar{K}-{\rm abs}}$ process is primarily due to the energy region being further separated from the reaction threshold.
@@
Figure 10: Fitting functions for (a1, a2, a3) the $\bar{K}\!N\!N$ production process and (b1, b2, b3) quasi-free kaon absorption for two-nucleon $QF_{\bar{K}-abs}$ process: (a1) and (b1) show spectral functions using the three-body phase space of $\Lambda pn$; (a2) and (b2) correspond to spectral functions using the four-body phase space of $\pi\Lambda NN$; (a3) and (b3) represent spectral functions using the four-body phase space of $\pi\Sigma pn$. The $z$ scale is logarithmic. The black, blue, and red dotted lines correspond to those in Fig. 6.
@@
This analysis assumes that the detected particles, $\pi Y\!N$, originate from the $(\bar{K}+NN)$ intermediate state. However, knocked-out $N^{\prime}$ in the primary $K^{-}N\to\bar{K}N^{\prime}$ reaction can be detected without detecting $N$ from the intermediate state. In such cases, the $(m,q)$ values for the $[\pi Y]N^{\prime}$ system (rather than the $[\pi Y\!N]$ system) are distributed widely across the kinematically allowed region. To incorporate this scenario, we generated expected spectral functions through simulations for each relevant process and summed them to account for the background, ${\rm BG}(m,~{}q,~{}m_{\pi Y})$. There is background contamination resulting from the misidentification of the final state, as described in Sec. III.2. In the present fitting, we omitted this background; however, we considered its effect when evaluating the cross-section systematic error.
@@
Because of the limited statistics for the limited acceptance on the three-dimensional kinematical space of $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]},\,m_{[\pi Y]})$, neither a multi-dimensional fitting nor the application of the coupled-channel-based sophisticated fitting function (to account for the threshold effects) is feasible in this analysis. Instead, we performed a simultaneous fit of three one-dimensional spectra along the $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$ axes. The two charged modes, $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$, were fitted simultaneously using common values of $M_{\Lambda(1405)}$ and $\Gamma_{\Lambda(1405)}$. The fitting results are shown in Fig. 11, and the obtained parameters are summarized in Table 1. The chi-square values and the number of degrees of freedom for $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ were $\chi^{2}/NDF=275.5/118$, $76.3/85$, and $114.9/136$, respectively. The figures demonstrate that the proposed model functions provide a good overall data description.
@@
Figure 11: Fitted distributions and results of fitting for channels (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. “BS”, “QF”, and “BG” denote $\bar{K}\!N\!N$ production, ${\rm QF}_{\bar{K}-{\rm abs}}$, and background processes, respectively.
Table 1: Parameters obtained by fitting.
@@
In Figs. 11(c) and 11(d), the $\Lambda(1405)$ contribution in the $m_{\pi^{\mp}\Sigma^{\pm}}$ distributions is well reproduced by a simple Breit–Wigner function. The obtained mass and width of $\Lambda(1405)$ are $1432\pm 5~{}{\rm MeV}/c^{2}$ and $49\pm 3~{}{\rm MeV}/c^{2}$, respectively. These values are close to the pole position deduced from the $d(K^{-},n)\pi\Sigma$ reactions, $(m,\Gamma)=(1418~{}{\rm MeV}/c^{2},52~{}{\rm MeV}/c^{2})$[18], and to the theoretically predicted pole positions based on chiral unitary approaches, $(1421-1434~{}{\rm MeV}/c^{2},20-52~{}{\rm MeV}/c^{2})$ [19, 20, 21, 22, 23, 24].
@@
We examined the validity of the global fit using $q_{[\pi Y\!N]}$-sliced mass spectra, as shown in Figs. 12 and 13. The model functions successfully reproduced the global structure of each sliced distribution except for the $m_{[\pi Y]}$ spectra near the $\bar{K}\!N$ threshold in the low-momentum-transfer region of $q<0.3~{}{\rm GeV}/c$, as shown in Fig.13. This discrepancy might arise from the effect of the neglected channel coupling with the $\bar{K}\!N$ channel (i.e., the $\bar{K}N$ threshold cusp in the $I_{\bar{K}N}=1$ channel recently reported by the Belle collaboration [25]).
@@
Figure 12: Differential cross-sections $d\sigma/dm_{\pi Y\!N}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fitting results obtained from the analysis. In panel (a), the gray hatched box highlights a region outside the fitting range to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N\!N}$ threshold.
@@
Figure 13: Differential cross-sections $d\sigma/dm_{\pi Y}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fitting results obtained from the analysis. In the case of (a), the requirements of $m_{\pi^{-}\Lambda p}<2.6~{}{\rm GeV}/c^{2}$ and $p_{\rm miss}>0.3~{}{\rm GeV}/c$ are imposed to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N}$ threshold.
@@
The total cross-sections of $\bar{K}\!N\!N$ decaying into each mesonic decay channel, $\sigma_{\bar{K}\!N\!N}^{tot}\,Br(\pi Y\!N)$, were evaluated by integrating the $\bar{K}\!N\!N$ fitting function over the entire kinematical region including above the $\bar{K}$ binding threshold, $m_{\bar{K}}+2m_{N}$, as summarized in Table 2. We also tabulated the branch below the $m_{\bar{K}}+2m_{N}$ threshold to exclude the energy region where the spectral shape of the fitting functions would become less reliable because of threshold effects. The values for non-mesonic decay modes, $\bar{K}\!N\!N_{I_{3}=+1/2}\to\Lambda p$ and $\Sigma^{0}p$, were taken from Ref. [2]. The values for the $\bar{K}\!N\!N\to\Sigma^{+}n$ and $\bar{K}\!N\!N\to\pi^{0}\Lambda p$ channels were evaluated as $Br(\Sigma^{+}n)=2Br(\Sigma^{0}p)$ and $Br(\pi^{0}\Lambda p)=1/2Br(\pi^{+}\Lambda n)$, respectively, on the basis of the isospin symmetry. The total non-mesonic decay branch was then evaluated as the sum of the three non-mesonic channels, the errors of which are the linear sum of these three channels because the errors are correlated.
@@
We found that the ratio of mesonic to non-mesonic decay branches is $\mathcal{O}(10)$ if we integrate all mass regions. The mesonic decay branch is still larger than the non-mesonic one, even if we integrate below the $\bar{K}$ binding threshold. These results suggest that the mesonic decay mode is the dominant decay branch of the $\bar{K}\!N\!N$, as expected from theoretical considerations.
@@
The decay branches $Br(\pi^{+}\Lambda n)$ and $Br(\pi^{\mp}\Sigma^{\pm}p)$ were similar in magnitude in both integration ranges. The results suggest that the mesonic decay modes coupled to the $I_{\bar{K}\!N}=1$ channel play a substantial role in the $\bar{K}\!N\!N$ decay process. In addition, this coupling could explain why the decay width of the $\bar{K}\!N\!N$ is much broader than that of the $\Lambda(1405)$ resonance: the mesonic decay modes with $I_{\bar{K}\!N}=1$ contribute to the broadening of the $\bar{K}\!N\!N$ decay width.
@@
Table 2: Cross-sections of each decay channel for $\bar{K}\!N\!N$. The first and second errors are statistical and systematic ones, respectively. The $\bar{K}$ binding thresholds for $I_{3}=\pm 1/2$ were considered as $m_{K^{-}}+2m_{p}$ and $m_{\bar{K}^{0}}+2m_{n}$, respectively. We estimated decay branches for $\Sigma^{+}n$ and $\pi^{0}\Lambda p$ channels by assuming isospin symmetry.
@@
Decay channel All regions Below the $\bar{K}$ binding threshold